General Products has an asset beta of 1.4, and a debt to equity ratio of 0.7. Their tax rate is 0.24. If the risk free rate is 0.02, and the expected return on the S&P500 is 0.13, what is the cost of levered equity for General Products?
Answer #1

Levered Beta = Asset Beta * [1 + {(D/E) * (1 - t)}]

= 1.4 * [1 + {0.7 * (1 - 0.24)}]

= 1.4 * [1 + 0.532] = 1.4 * 1.532 = 2.1448

According to the CAPM,

Cost of Levered Equity = Risk-free Rate + [Levered Beta * (Expected Market Return - Risk-free Rate)]

= 0.02 + [2.1448 * (0.13 - 0.02)] = 0.02 + [2.1448 * 0.11] = 0.02 + 0.2359 = 0.2559, or 25.59%